# Optimal Average Disk-Inspection via Fermat’s Principle

### v2

@@ Abstract
This work resolves the optimal average-case cost of the Disk-Inspection problem, a variant of Bellman’s 1955 lost-in-a-forest problem. In Disk-Inspection, a mobile agent starts at the center of a unit disk and follows a trajectory that inspects perimeter points whenever the disk does not obstruct visibility. The worst-case cost was solved optimally in 1957 by Isbell [30], but the average-case version remained open, with heuristic upper bounds proposed by Gluss [28] in 1961 and improved only recently in [16].
Our approach applies Fermat’s Principle of Least Time to the discretization framework of [16], showing that optimal solutions are captured by a one-parameter family of recurrences independent of the discretization size. In the continuum limit these recurrences give rise to a single-parameter optimal control problem, whose trajectories coincide with limiting solutions of the original Disk-Inspection problem. A crucial step is proving that the optimal initial condition generates a trajectory that avoids the unit disk, thereby validating the optics formulation and reducing the many-variable optimization to a rigorous one-parameter problem. In particular, this disproves Gluss’s conjecture [28] that optimal trajectories must touch the disk.
@@ 1 Introduction
The above setting is Bellman’s 1955 lost-in-a-forest problem specialized to a halfspace with known distance, commonly called the shoreline problem with known distance. Isbell [30] solved the worst-case version optimally in 1957. The average-case variant was first examined by Gluss [28], who proposed heuristic search strategies to the problem. He further conjectured that optimal solutions, as in the worst-case setting, must touch the unit disk. More recently, the same problem was studied again from the equivalent perspective of Disk-Inspection, leading to a high-dimensional nonconvex NonLinear Program (NLP) [16] modelling a discretized version of the problem, which achieved an average-cost upper bound of $3.5509015$. However, that approach could not certify global optimality of the NLP solutions (which would imply that the reported bound is close to the true optimum), the derived solutions did not scale computationally (which would theoretically give better upper bounds), and it offered no structural characterization of optimal trajectories in the continuous setting. Moreover, no nontrivial lower bound for the average-case cost was known, leaving it entirely unclear how close the reported bound was to the truth.
In this work, we resolve the Average-Case Disk-Inspection problem. We prove that the exact optimum is $3.54925958\ldots,$ accurate to at least six digits, and that it is realized by trajectories determined through the solutions of an ODE system. Our key technical contribution is to reformulate the discretized problem, via Fermat’s Principle of Least Time, as an optics problem. Assuming the optimal trajectory does not touch the disk, we show that optimal solutions to the discrete inspection problem admit a recurrence structure that, in the continuum limit, yields a single-parameter optimal control problem. In this optimization problem, feasible solutions are trajectories of an ODE system determined by one initial condition, and optimizing over this parameter gives a rigorously certifiable optimum. A crucial part of the analysis is to prove that the non-touching condition is in fact satisfied, so the reduction to the optics problem is valid. Overall, this transforms the previous nonconvex many-variable optimization of [16] with no guarantees into a tractable one-parameter problem, completing the proof of optimality and establishing that the optimal trajectory avoids the unit disk, contrary to Gluss’s conjecture [28], thereby settling the optimal solution to the average-case problem definitively.
@@ 1.1 Related Work
In the mid-50s, Bellman [9] introduced what is nowadays called lost-in-a-forest-problem, one of the earliest formal questions in search theory. The input is a region $R$ (the forest) containing a point $P$ (starting position of a searcher/mobile agent) chosen at random, and the objective is to design a trajectory that minimizes the expected time to reach the boundary, starting from $P$. This formulation became a cornerstone for later work on search under uncertainty, and still many variations that have been proposed through the decades remain open.
@@
Surveys of the lost-in-a-forest problem are given in [10, 22]. Work on specific domains includes convex polygons [27], strips [22] and general regions analyzed through competitive ratios [33]. When region $R$ is a half-space, and the starting position of the agent is fixed deterministically, the problem is known as the shoreline problem and was first studied for a single agent in [7] and later extended to several agents in [5, 6, 8, 31]. For one agent, the logarithmic spiral is the best strategy known, with ratio $13.81$ [7, 23]. The strongest lower bounds known are $6.3972$ in general [6] and $12.5385$ for cyclic searches [35]. For two agents, a double logarithmic spiral achieves ratio $5.2644$ [6]. The optimal solutions to the one and two agent search problems are still open. For $n\geq 3$, trajectories along rays achieve at most $1/\cos(\pi/n)$ [6], and lower bounds matching these values were proven for $n\geq 4$ and for $n=3$ in [1] and [18], respectively.
@@
The variation to lost-in-a-forest pertaining to our results is the shoreline problem with known distance where the searcher knows her distance to the boundary of the forest. For minimizing the worst-case cost, the problem was solved optimally by Isbell [30] for one agent, and by Dobrev et al. in [18] for two agents, while [16] recently extended the results to the case where the hidden shorelines are tangent to a contiguous portion of the disk, rather than the entire disk.
@@
The average-case version of the shoreline problem with known distance was first studied by Gluss [28], who proposed two heuristic strategies and carried out a rigorous expected-cost analysis for each. He further conjectured that the optimal trajectory, as in the worst-case setting, must touch the disk. The numerical values reported in [28] were miscalculated, and as explained in the full version [15] of the conference paper [16], the best heuristic bound due to Gluss is $3.63489\ldots$. The first systematic approach to the average-case problem was given in the conference version [16] and its full version [15]. Their method introduced a discretization framework and formulated a nonconvex NonLinear Program (NLP) whose feasible solutions correspond to valid inspection trajectories, with the objective exactly capturing their average cost. Solving this NLP for large discretization parameter $k$ produced a sequence of feasible trajectories and rigorous upper bounds, resulting in a reported value of $3.5509015\ldots$. However, because the NLP is nonconvex, the obtained solutions could not be certified as globally optimal, and its computational hardness restricted $k$ to moderate values, preventing sharper approximations. No lower bounds for the average-case cost had been studied or reported, so it remained unclear how close the best previously reported bound was to the true optimum.
@@
Beyond shoreline search and lost-in-a-forest formulations, related problems in search theory have been studied extensively. The field itself is well established, with books and surveys providing systematic accounts of its models and applications [2, 3, 4, 17, 25]. Within point search in the plane, Langetepe [34] proved that the logarithmic spiral is optimal for a single robot, while [26] recently analyzed the case of multi-speed agents. Other studies considered search for a circle in a plane [29], multi-agent and grid searches with memory constraints [20, 21, 24, 36], and more recent work investigated searches in geometric terrains [12, 37] and the role of information in determining search cost [38, 39], to name only a few.
@@ 2.1 Problem Definition and the Main Result
The Worst-Case Disk-Inspection problem was solved in [30], with optimal cost $1+\sqrt{3}+7\pi/6\approx 6.39724$. Inspective trajectories admit two equivalent descriptions, see [16]. First, they inspect every point of the unit disk in the following sense. For any perimeter point $P$ there exists a point $A$ on the trajectory such that $\|\lambda A+(1-\lambda)P\|\geq 1$ for all $\lambda\in[0,1]$ (that is, $A$ sees $P$ without obstruction by the disk). Second, the trajectories intersect every line tangent to the unit disk (i.e. they discover eventually all shorelines). In what follows we use the inspection perspective, which will also define the average-case problem.
@@
The average-case problem was first studied by Gluss [28], who proposed two heuristic strategies and conjectured that optimal trajectories touch the disk. More recently, a discretization-based nonlinear programming framework was developed [15, 16], yielding the first systematic upper bounds, with best reported value $3.5509015$. However, the nonconvexity of this approach precluded global optimality guarantees, and no lower bounds were known, so the exact optimum remained unresolved. We resolve this problem with the following result, establishing the optimal value of ADI with high numerical accuracy. In particular, we prove that the previously reported upper bound in [16] was indeed very close to optimal, and we obtain this conclusion through a new continuous framework that replaces the earlier discrete, upper-bound approach. The technical contributions underlying this framework are presented in Section 2.2. We now state our main result.
@@
The optimal solution to ADI has cost $3.549259\ldots$.
@@ 2.2 New Technical Contributions
Our second contribution is a continuous framework that replaces the discrete approach of [16]. In this framework, candidate inspection trajectories are described, partially, by trajectories generated from an ordinary differential equation (ODE) system. The system is governed by a single real parameter, which turns ADI into a one-parameter optimal control problem. We now introduce the ODE system and the associated trajectories, which form the basis of the technical result leading to Theorem 2.4.
@@
It is known that an optimal trajectory to ADI starts with a so-called deployment phase. This is a line segment connecting the center of the disk to some point $A_{\infty}$ on the line $x=1$, where the slope $\theta$ of segment $OA_{\infty}$ will be referred to as the deployment angle. The reader may consult Figure 1 which shows the optimal inspective trajectory certifying Theorem 2.4. The initial deployment phase is followed by the so-called inspection phase shown as the green curve whose other endpoint is $A_{0}$. In the discrete setting (Figure 3), the endpoint is denoted $A_{k}$, where $k$ is proportional to a discretization parameter that tends to infinity. It is therefore natural to denote the limiting endpoint by $A_{\infty}$.
Figure 1: The optimal inspective trajectory certifying Theorem 2.4. The trajectory starts with the deployment phase $OA_{\infty}$, and is followed by the green curve (inspection phase) $\mathcal{T}^{\tau_{0}}$, solution to $\textsc{Sys}(\tau_{0})$, for $\tau_{0}\approx 1.64697686$, which is inspection-feasible with deployment parameter $\xi(\tau_{0})\approx 0.821619$. The inspection phase is therefore the curve $\mathcal{T}^{\tau_{0}}(x)$ for $x\in[0,\xi(\tau_{0})]$, with endpoints $A_{0}=\mathcal{T}^{\tau_{0}}(0)$ and $A_{\infty}=\mathcal{T}^{\tau_{0}}(\xi(\tau_{0}))$. The dotted blue line marks $x=1$, corresponding to the feasibility condition $\mathcal{T}^{\tau_{0}}_{1}(\xi)=1$. The deployment angle is $\theta=(1-\xi(\tau_{0}))\pi$, with labelled points $P_{0}=(1,0)$, $A_{\infty}=(1,\tan\mkern-2.0mu\left(\theta\right))$, and $A_{0}=(1,-\tau_{0})$.
@@
If the inspection phase of the optimal trajectory for ADI does not touch the unit disk, then the inspection phase is given as the ODE-generated curve $\mathcal{T}^{\tau_{0}}$ for some inspection‑feasible $\tau_{0}$, where $(\psi,\tau)$ solves $\textsc{Sys}(\tau_{0})$, and $\xi=\xi(\tau_{0})>1/2$ is the deployment parameter of $\tau_{0}$. Together with the deployment phase $O\mathcal{T}^{\tau_{0}}(\xi)$, the optimal cost to ADI is
@@
minimized over all inspection‑feasible $\tau_{0}\in\mathcal{I}$, where $\theta=(1-\xi)\pi$ is the corresponding deployment angle.
@@ 3 Background Machinery
We begin by reviewing the main tools introduced in [16], which we follow throughout Section 3 apart from minor reformulations. These ideas play only a preliminary role in our development, since they provide the initial reduction and notation on which our main arguments build. That earlier work proposed a discretized version of ADI together with a Nonlinear Programming (NLP) formulation whose solutions yield upper bounds. In Section 3.1 we introduce a related intermediate “partial” problem that simplifies the reduction from the continuous setting. Section 3.2 then describes the discretized problem and concludes with a brief sketch of the NLP formulation, which, while not directly relevant here, was previously used to derive upper bounds to ADI.
@@ 3.1 Reduction to the Partial Average-Case Disk-Inspection Problem
Any $\theta$-inspective curve can be extended to an inspective trajectory by appending the line segment from the origin to $(1,\tan\mkern-2.0mu\left(\theta\right))$ (previously referred to as the deployment phase). One of the results in [16] was the following explicit relation between the two problems.
@@
The task is therefore reduced to selecting the deployment angle $\theta$ and designing the corresponding $\theta$-inspective curve of minimal average cost $s_{0}(\theta)$. The starting point for our work is the formulation that lead to the upper bound on $B_{\theta}(s_{0})$ established in [16].

### v3

# Konstantinos Georgiou

@@ Abstract
This work resolves the optimal average-case cost of the Disk-Inspection problem, a variant of Bellman’s 1955 lost-in-a-forest problem. In Disk-Inspection, a mobile agent starts at the center of a unit disk and follows a trajectory that inspects perimeter points whenever the disk does not obstruct visibility. The worst-case cost was solved optimally in 1957 by Isbell [isbell1957optimal], but the average-case version remained open, with heuristic upper bounds proposed by Gluss [gluss1961alternative] in 1961 and improved only recently in [conley2025multi].
Our approach applies Fermat’s Principle of Least Time to the discretization framework of [conley2025multi], showing that optimal solutions are captured by a one-parameter family of recurrences independent of the discretization size. In the continuum limit these recurrences give rise to a single-parameter optimal control problem, whose trajectories coincide with limiting solutions of the original Disk-Inspection problem. A crucial step is proving that the optimal initial condition generates a trajectory that avoids the unit disk, thereby validating the optics formulation and reducing the many-variable optimization to a rigorous one-parameter problem. In particular, this disproves Gluss’s conjecture [gluss1961alternative] that optimal trajectories must touch the disk.
@@ 1 Introduction
The above setting is Bellman’s 1955 lost-in-a-forest problem specialized to a halfspace with known distance, commonly called the shoreline problem with known distance. Isbell [isbell1957optimal] solved the worst-case version optimally in 1957. The average-case variant was first examined by Gluss [gluss1961alternative], who proposed heuristic search strategies to the problem. He further conjectured that optimal solutions, as in the worst-case setting, must touch the unit disk. More recently, the same problem was studied again from the equivalent perspective of Disk-Inspection, leading to a high-dimensional nonconvex NonLinear Program (NLP) [conley2025multi] modelling a discretized version of the problem, which achieved an average-cost upper bound of $3.5509015$. However, that approach could not certify global optimality of the NLP solutions (which would imply that the reported bound is close to the true optimum), the derived solutions did not scale computationally (which would theoretically give better upper bounds), and it offered no structural characterization of optimal trajectories in the continuous setting. Moreover, no nontrivial lower bound for the average-case cost was known, leaving it entirely unclear how close the reported bound was to the truth.
In this work, we resolve the Average-Case Disk-Inspection problem. We prove that the exact optimum is $3.54925958\ldots,$ accurate to at least six digits, and that it is realized by trajectories determined through the solutions of an ODE system. Our key technical contribution is to reformulate the discretized problem, via Fermat’s Principle of Least Time, as an optics problem. Assuming an optimal trajectory does not touch the disk, we show that optimal solutions to the discrete inspection problem admit a recurrence structure that, in the continuum limit, yields a single-parameter optimal control problem. In this optimization problem, feasible solutions are trajectories of an ODE system determined by one initial condition, and optimizing over this parameter gives a rigorously certifiable optimum. A crucial part of the analysis is to prove that the non-touching condition is in fact satisfied, so the reduction to the optics problem is valid. Overall, this transforms the previous nonconvex many-variable optimization of [conley2025multi] with no guarantees into a tractable one-parameter problem, completing the proof of optimality and establishing that an optimal trajectory avoids the unit disk, contrary to Gluss’s conjecture [gluss1961alternative], thereby settling the optimal solution to the average-case problem definitively.
@@ 1.1 Related Work
In the mid-50s, Bellman [bellman1956minimization] introduced what is nowadays called lost-in-a-forest-problem, one of the earliest formal questions in search theory. The input is a region $R$ (the forest) containing a point $P$ (starting position of a searcher/mobile agent) chosen at random, and the objective is to design a trajectory that minimizes the expected time to reach the boundary, starting from $P$. This formulation became a cornerstone for later work on search under uncertainty, and still many variations that have been proposed through the decades remain open.
@@
Surveys of the lost-in-a-forest problem are given in [berzsenyi1995lost, finch2004lost]. Work on specific domains includes convex polygons [gibbs2016bellman], strips [finch2004lost] and general regions analyzed through competitive ratios [kubel2021approximation]. When region $R$ is a half-space, and the starting position of the agent is fixed deterministically, the problem is known as the shoreline problem and was first studied for a single agent in [baeza1988searching] and later extended to several agents in [baeza1997searching, baeza1995parallel, baezayates1993searching, jez2009two]. For one agent, the logarithmic spiral is the best strategy known, with ratio $13.81$ [baeza1988searching, finch2005searching]. The strongest lower bounds known are $6.3972$ in general [baeza1995parallel] and $12.5385$ for cyclic searches [langetepe2012searching]. For two agents, a double logarithmic spiral achieves ratio $5.2644$ [baeza1995parallel]. The optimal solutions to the one and two agent search problems are still open. For $n\geq 3$, trajectories along rays achieve at most $1/\cos(\pi/n)$ [baeza1995parallel], and lower bounds matching these values were proven for $n\geq 4$ and for $n=3$ in [AcharjeeGKS19] and [dobrev2020improved], respectively.
@@
The variation to lost-in-a-forest pertaining to our results is the shoreline problem with known distance where the searcher knows her distance to the boundary of the forest. For minimizing the worst-case cost, the problem was solved optimally by Isbell [isbell1957optimal] for one agent, and by Dobrev et al. in [dobrev2020improved] for two agents, while [conley2025multi] recently extended the results to the case where the hidden shorelines are tangent to a contiguous portion of the disk, rather than the entire disk.
@@
The average-case version of the shoreline problem with known distance was first studied by Gluss [gluss1961alternative], who proposed two heuristic strategies and carried out a rigorous expected-cost analysis for each. He further conjectured that any optimal trajectory, as in the worst-case setting, must touch the disk. The numerical values reported in [gluss1961alternative] were miscalculated, and as explained in the full version [conley2025multiagentdiskinspectionArxiv] of the conference paper [conley2025multi], the best heuristic bound due to Gluss is $3.63489\ldots$. The first systematic approach to the average-case problem was given in the conference version [conley2025multi] and its full version [conley2025multiagentdiskinspectionArxiv]. Their method introduced a discretization framework and formulated a nonconvex NonLinear Program (NLP) whose feasible solutions correspond to valid inspection trajectories, with the objective exactly capturing their average cost. Solving this NLP for large discretization parameter $k$ produced a sequence of feasible trajectories and rigorous upper bounds, resulting in a reported value of $3.5509015\ldots$. However, because the NLP is nonconvex, the obtained solutions could not be certified as globally optimal, and its computational hardness restricted $k$ to moderate values, preventing sharper approximations. No lower bounds for the average-case cost had been studied or reported, so it remained unclear how close the best previously reported bound was to the true optimum.
@@
Beyond shoreline search and lost-in-a-forest formulations, related problems in search theory have been studied extensively. The field itself is well established, with books and surveys providing systematic accounts of its models and applications [ahlswede1987search, alpern2013search, alpern2006theory, CGK19search, gal2010search]. Within point search in the plane, Langetepe [langetepe2010optimality] proved that the logarithmic spiral is optimal for a single robot, while [georgiou2025spiralsLATIN] recently analyzed the case of multi-speed agents. Other studies considered search for a circle in a plane [gluss1961minimax], multi-agent and grid searches with memory constraints [emek2015many, Emekicalp2014, fricke2016distributed, LangnerKUW15], and more recent work investigated searches in geometric terrains [bouchard2018deterministic, pelc2018reaching] and the role of information in determining search cost [pelc2018information, pelc2019cost], to name only a few.
@@ 2.1 Problem Definition and the Main Result
The Worst-Case Disk-Inspection problem was solved in [isbell1957optimal], with optimal cost $1+\sqrt{3}+7\pi/6\approx 6.39724$. Inspective trajectories admit two equivalent descriptions, see [conley2025multi]. First, they inspect every point of the unit disk in the following sense. For any perimeter point $P$ there exists a point $A$ on the trajectory such that $\|\lambda A+(1-\lambda)P\|\geq 1$ for all $\lambda\in[0,1]$ (that is, $A$ sees $P$ without obstruction by the disk). Second, the trajectories intersect every line tangent to the unit disk (i.e. they discover eventually all shorelines). In what follows we use the inspection perspective, which will also define the average-case problem.
@@
The average-case problem was first studied by Gluss [gluss1961alternative], who proposed two heuristic strategies and conjectured that optimal trajectories touch the disk. More recently, a discretization-based nonlinear programming framework was developed [conley2025multiagentdiskinspectionArxiv, conley2025multi], yielding the first systematic upper bounds, with best reported value $3.5509015$. However, the nonconvexity of this approach precluded global optimality guarantees, and no lower bounds were known, so the exact optimum remained unresolved. We resolve this problem with the following result, establishing the optimal value of ADI with high numerical accuracy. In particular, we prove that the previously reported upper bound in [conley2025multi] was indeed very close to optimal, and we obtain this conclusion through a new continuous framework that replaces the earlier discrete, upper-bound approach. The technical contributions underlying this framework are presented in Section 2.2. We now state our main result.
@@
$\lvert C_{\mathrm{opt}}-3.549259\rvert<10^{-7}$, where $C_{\mathrm{opt}}$ is the cost for an optimal solution to ADI.
@@ 2.2 New Technical Contributions
Our second contribution is a continuous framework that replaces the discrete approach of [conley2025multi]. In this framework, candidate inspection trajectories are described, partially, by trajectories generated from an ordinary differential equation (ODE) system. The system is governed by a single real parameter, which turns ADI into a one-parameter optimal control problem. We now introduce the ODE system and the associated trajectories, which form the basis of the technical result leading to Theorem 2.4.
@@
Following the standard shoreline framework (see, e.g., [conley2025multi], building on [gluss1961alternative]), we restrict attention to ADI trajectories that start with a a so-called deployment phase. This is defined as a line segment connecting the center of the disk to some point $A_{\infty}$ on the line $x=1$, where the slope $\theta$ of segment $OA_{\infty}$ will be referred to as the deployment angle, and hence $A_{\infty}=(1,\tan\theta)$ with $\theta\in[0,\pi/2)$. The reader may consult Figure 1 which shows the optimal inspective trajectory certifying Theorem 2.4. The initial deployment phase is followed by the so-called inspection phase shown as the green curve whose other endpoint is $A_{0}$. In the discrete setting (Figure 3), the endpoint is denoted $A_{k}$, where $k$ is proportional to a discretization parameter that tends to infinity. It is therefore natural to denote the limiting endpoint by $A_{\infty}$.
Figure 1: The optimal inspective trajectory certifying Theorem 2.4. The trajectory starts with the deployment phase $OA_{\infty}$, and is followed by the green curve (inspection phase) $\mathcal{T}^{\tau_{0}}$, solution to $\textsc{Sys}(\tau_{0})$, for $\tau_{0}\approx 1.64697686$, which is inspection-feasible with deployment parameter $\xi(\tau_{0})\approx 0.81190987$. The inspection phase is therefore the curve $\mathcal{T}^{\tau_{0}}(x)$ for $x\in[0,\xi(\tau_{0})]$, with endpoints $A_{0}=\mathcal{T}^{\tau_{0}}(0)$ and $A_{\infty}=\mathcal{T}^{\tau_{0}}(\xi(\tau_{0}))$. The dotted blue line marks $x=1$, corresponding to the feasibility condition $\mathcal{T}^{\tau_{0}}_{1}(\xi)=1$. The deployment angle is $\theta=(1-\xi(\tau_{0}))\pi$, with labelled points $P_{0}=(1,0)$, $A_{\infty}=(1,\tan\mkern-2.0mu\left(\theta\right))$, and $A_{0}=(1,-\tau_{0})$.
@@
If the inspection phase of an optimal trajectory for ADI does not touch the unit disk, then the inspection phase is the subcurve $\left(\mathcal{T}^{\tau_{0}}(x)\right)_{0\leq x\leq\xi}$ for some inspection-feasible $\tau_{0}$, where $(\psi,\tau)$ solves $\textsc{Sys}(\tau_{0})$, and $\xi=\xi(\tau_{0})>1/2$ is the deployment parameter of $\tau_{0}$. Together with the deployment segment $O\mathcal{T}^{\tau_{0}}(\xi)$, the optimal cost to ADI is
@@
minimized over all inspection-feasible $\tau_{0}\in\mathcal{I}$, where $\theta=(1-\xi)\pi$ is the corresponding deployment angle.
@@ 3 Background Machinery
We begin by reviewing the main tools introduced in [conley2025multi], which we follow throughout Section 3 apart from minor reformulations. These ideas play only a preliminary role in our development, since they provide the initial reduction and notation on which our main arguments build. That earlier work proposed a discretized version of ADI together with a Nonlinear Programming (NLP) formulation whose solutions yield upper bounds. In Section 3.1 we introduce a related intermediate “partial” problem that simplifies the reduction from the continuous setting. Section 3.2 then describes the discretized problem and concludes with a brief sketch of the NLP formulation, which, while not directly relevant here, was previously used to derive upper bounds to ADI.
@@ 3.1 Reduction to the Partial Average-Case Disk-Inspection Problem
Any $\theta$-inspective curve can be extended to an inspective trajectory by appending the line segment from the origin to $(1,\tan\mkern-2.0mu\left(\theta\right))$ (previously referred to as the deployment phase). One of the results in [conley2025multi] was the following explicit relation between the two problems.
@@
The task is therefore reduced to selecting the deployment angle $\theta$ and designing the corresponding $\theta$-inspective curve of minimal average cost $s_{0}(\theta)$. The starting point for our work is the formulation that lead to the upper bound on $B_{\theta}(s_{0})$ established in [conley2025multi].
Theorem 3.3 is the continuous analogue of Lemma 3 in [conley2025multi], which is stated for a discrete inspection setting. The proof there is geometric and does not rely on discreteness: the total average cost splits into two independent contributions, namely (i) the average cost of inspecting the initial arc of length $2\theta$ during the deployment phase, and (ii) the average cost of inspecting the remaining portion of the perimeter, regardless of whether it consists of finitely many or infinitely many points.
Every feasible ADI trajectory admits such a decomposition. Following the classical shoreline analysis of Isbell and Gluss, one may restrict attention to trajectories whose rim begins with a straight deployment segment from the origin to a point $(1,\tan\theta)$ for some $\theta\in[0,\pi/2)$, and is angularly monotone thereafter. For a fixed $\theta$, the contribution of the continuation to the total ADI cost depends only on the average cost of the corresponding $\theta$-ADI solution. If the continuation were not optimal for that $\theta$-ADI instance, replacing it by a better $\theta$-ADI curve would strictly improve the total ADI cost. Hence, the continuation of any optimal ADI solution must itself be an optimal solution of the associated $\theta$-ADI problem.